Question Help In a survey of 3297 adults aged 57 through 85 years, it was found that 80.3% of the a. How many of the 3297 subjects used at least one prescription medication? used at least one prescription medication. Complete parts (a) through Round to the nearest integer as needed.) b. Construct a 90% confidence interval estimate of the percentage of adults aged 57 through 85 years who use at least one prescription medication. Round to one decimal place as needed.) c. What do the results tell us about the proportion of college students who use at least one prescription medication? The results tell us nothing about the proportion of college students who use at least one prescription medication The results tell us that, with 9 % confidence, the true proportion of college students who use at least one prescription medicati n is it e inte o A. О В al found in part The results tell us that, with 9 % confidence, the probability that a college student uses at least one prescription medication is in the ite vah The results tell us that there is a 90% probability that the true proportion of college students who use at east one rescript on medic son h he in part (b) С D. n at b e nd
Answer #1

a)

n = 3297

p = 0.803

No. of students = 2648

b)

Standard error of the mean = SEM = √x(N-x)/N3 = 0.007

α = (1-CL)/2 = 0.050

Standard normal deviate for α = Zα = 1.645

Proportion of positive results = P = x/N = 0.803

Lower bound = P - (Zα*SEM) = 0.80

Upper bound = P + (Zα*SEM) = 0.8

c) B option